# Preference Poisoning Attacks on Reward Model Learning

### v1

@@ 4. Threat Model
We consider the presence of a malicious actor (adversary) who can modify preference labels for pairs of outcomes in the dataset. Since these are inherently subject to human feedback, with no objective way to ground them (being expressions of human preferences), it is not possible to verify responses received against some ground truth. Moreover, preferences of this kind are commonly obtained over the internet, the relative anonymity of which provides ample opportunity for malicious subversion of preference information. In particular, attackers can enter as annotators of the system, solicited, for example, using Amazon Mechanical Turk, Prolific, etc, or hired using other means (e.g., by using a third-party outsourcing service, such as Sama (Perrigo, 2023)), or can pay off a subset of such annotators.
4.1. Attacker Capabilities and Constraints
@@
Formally, we assume that the attacker can modify (flip) a subset of preference labels $o_{i}$ for corresponding pairs of input outcomes $(x_{i},y_{i})$, but not the feature vectors associated with these outcomes (as these are typically pre-specified as part of the preference elicitation process). Since preferences are commonly elicited from a large number of users as a way to reliably learn reward models, we assume that the adversary can only interject malicious preferences over a subset of these. For example, if this is done through malware on a user’s device that submits fake pairwise comparison results on the user’s behalf, such attacks can affect only a subset of users. Similarly, if malicious users interject themselves into the elicitation process (something that is difficult to prevent at scale, given the relative anonymity inherent in it), there is a limit to how many such malicious accounts can be created (or a limit on the cost an attacker can expend on using botnets for such a task). Finally, if attackers simply pay off regular hired annotators to submit malicious preferences, there is a limit on how much the adversary can spend.
@@
Formally, we allow at most $B$ labels to be flipped. Let $\tilde{D}=\{(x_{i},y_{i},\tilde{o}_{i})\}$ be the dataset produced after the attacker manipulates (up to) $B$ labels in $D$, and let $\mathcal{L}(\tilde{D},\theta)$ be the loss function optimized by the learner over the poisoned dataset $\tilde{D}$. Let $\delta_{i}\in\{-1,0,+1\}$ denote the label perturbations by the attacker, so that $\tilde{o}_{i}=o_{i}+\delta$. The constraint is then $\sum_{i\in D}|\delta_{i}|\leq B$. We make explicit the dependence of poisoned data on $\delta$ using the notation $\tilde{D}(\delta)$.
4.2. Attacker Goals
@@
A challenge in formalizing a promotion attack is to make precise what “others” would be with which outcomes in $C^{T}$ will be compared. We propose to model this by supposing an unknown distribution $\mathcal{P}$ over outcomes, with the adversarial goal being to maximize induced average success rate of outcomes in $C^{T}$ over a randomly drawn outcome $x\sim\mathcal{P}$ (that is, $x$ drawn according to $\mathcal{P}$). The distribution $\mathcal{P}$ captures an exogenous distribution over outcomes (e.g., products, prompt responses, trajectories) that will be generated in whatever context is relevant, and with which the target outcomes will be compared. We therefore model the promotion attack as solving the following optimization problem:
@@
(3) $\begin{split}&\max_{\delta}F(\delta)\equiv\sum_{c\in C^{T}}\Pr_{x\sim\mathcal{P}}\{R_{\theta}(c)\geq R_{\theta}(x)\}\\ &\mathrm{s.t.:}\quad\theta\in\arg\max_{\theta^{\prime}}\mathcal{L}(\tilde{D}(\delta),\theta^{\prime}).\end{split}$
@@
An evident technical challenge with this attack model is that it is impossible to evaluate $\Pr_{x\sim\mathcal{P}}\{R_{\theta}(c)\geq R_{\theta}(x)\}$ since $\mathcal{P}$ is unknown. Moreover, even if we knew $\mathcal{P}$, computing this probability exactly can be computationally intractable. Practically, therefore, we would instead estimate it using a collection of sample outcomes $\{x_{i}\}_{i=1}^{N}$, which we assume are drawn from the unknown distribution $\mathcal{P}$ (a conventional assumption in learning theory, for example (Anthony and Bartlett, 1999)). In practice, we would draw them, for example, from the dataset we aim to poison. This yields the following finite-sample approximation of Problem (3):
@@
(4) $\begin{split}&\max_{\delta}\hat{F}(\delta)\equiv\sum_{c\in C^{T}}\sum_{i}\mathbf{1}_{R_{\theta}(c)\geq R_{\theta}(x_{i})}\\ &\mathrm{s.t.:}\quad\theta\in\arg\max_{\theta^{\prime}}\mathcal{L}(\tilde{D}(\delta),\theta^{\prime}),\end{split}$
where $\mathbf{1}_{R_{\theta}(c)\geq R_{\theta}(x_{i})}$ is 1 whenever the condition is true, and 0 otherwise. A natural question is whether this approximation yields a solution that approximates the true optimization problem that the attacker aims to solve—that is, Problem (3). Note that this is not self-evident. First, we are not merely approximating the expectation with sample average (and removing the constant factor, which does not impact the optimal solution), but approximating an optimization problem. Thus, law of large numbers does not immediately imply that the quality of solution to the approximate problem converges. Second, the problem involves bi-level optimization, since the attacker’s optimal solution depends, in turn, on the learning optimization problem that yields the reward model parameters $\theta$. Next, we prove that this problem has polynomial sample complexity, that is, the number of samples $N$ grows only polynomially in $1/\epsilon$ and the dimension of the function class $R_{\theta}$, where $\epsilon$ represents the quality of the approximation of the solution to Problem (3) by the solution to Problem (4).
Lemma 4.1 ().
[…]
Proof.
Theorem 4.2 (Anthony and Bartlett (1999)).
[…]
Theorem 4.3 ().
[…]
Proof.
Similarly, Problem (4) is equivalent (up to the $1/N$ factor) to
Demotion Attacks
@@
Demotion attacks are a mirror image of promotion attacks: the attacker goal in this case is to demote a target set of candidates $C^{T}$ in terms of the comparisons with others induced by the learned reward model $R_{\theta}$. For example, a firm in a concentrated market would wish to prevent would want to demote the perceived preference of its competitor’s product, or a malicious actor would wish to prevent certain useful information from being shown by conversational AI (e.g., vaccine information in response to prompts about an infectious disease). We formalize it in precisely the same manner as we did for promotion attacks, with the goal now to minimize $F(\delta)$, rather maximize it. We can again use $\hat{F}(\delta)$ as the approximate objective to this end. The result in Theorem 4.3 can be directly adapted to this setting as well.
4.3. Attacker Information
We assume that the attacker knows the dataset $D$ that they are modifying (or whatever subset of $D$ is available to the attacker for possible modification of pairwise comparison labels). We consider two information settings vis-a-vis the model architecture itself: white-box and black-box. In the white-box setting, the attacker is assumed to know the true model (e.g., neural network) architecture being trained. In the black-box setting, the attacker does not know the architecture. Notably, we do not assume knowledge of the random initialization in training.
4.4. Stealth
@@
An important consideration in poisoning attacks is stealth, that is, ensuring that the fact of the attack is not immediately evident to observers. To this end, we use degradation in overall test accuracy as a measure of stealth: we view the attack as stealthy if test accuracy does not significantly degrade. We treat stealth as a post-hoc requirement on the attack, that is, we experimentally check that the attack does not have a significant accuracy degradation as a side-effect.
5. Attack Algorithms
We develop two classes of algorithmic approaches for approximately identifying which preference comparisons to flip. The first class of approaches are based on gradient-based techniques. The key challenges in using gradient-based methods in our setting are (a) the bi-level nature of the optimization problem, (b) the fact that the problem is discrete (flipping labels), and (c) the discontinuous and non-differentiable nature of the objective. The second class are greedy rank-by-distance approaches, in which the main variation is how we measure distance. We detail these next.
5.1. Gradient-Based Framework
@@
(5) $U(\delta)=\sum_{c\in C^{T}}R_{\theta}(c)-\frac{|C^{T}|}{N}\sum_{i}R_{\theta}(x_{i}).$
Now, since the only dependence on the decision variables $\delta$ is via the parameters learned $\theta$, as long as $R_{\theta}$ are differentiable with respect to $\theta$, the objective will pose no issues. Next, we turn to the problem of discrete decision space $\delta$, where $\delta_{i}\in\{-1,0,+1\}$. We propose to relax $\delta_{i}\in[-1,+1]$, and use projection after a series of gradient steps to subsequently output a feasible $\delta$ vector. The projection step will choose the $B$ $\delta_{i}$s with the largest magnitude as the attack, setting the rest to zero, and then round the selected $\delta_{i}$s.
@@
where we now make explicit the dependence of $\theta$ on $\delta$, and $\frac{d\theta(\delta)}{d\delta}$ is the matrix of derivatives $\frac{d\theta_{j}}{d\delta_{i}}$. While computing $\nabla_{\theta}R_{\theta}$ is straightforward, $\theta(\delta)$ is an implicit function of $\delta$ that arises due to the indirect dependence of learned parameters $\theta$ on $\delta$ as solutions to the optimization problem $\min_{\theta^{\prime}}\mathcal{L}(\tilde{D}(\delta),\theta^{\prime})$. Computing the implicit derivatives $\frac{d\theta(\delta)}{d\delta}$ is a common issue in gradient-based methods for poisoning attacks (see, e.g., Mei and Zhu (2015)), and a common approach is to leverage the implicit function theorem (Koh and Liang, 2017; Mei and Zhu, 2015; Simon et al., 1994). Specifically, since MLE in our context is an unconstrained optimization problem, an optimal solution satisfies first-order conditions $\nabla_{\theta}\mathcal{L}(\tilde{D}(\delta),\theta)=0$. With mild assumptions detailed elsewhere (Simon et al., 1994), we compute the implicit derivative we need as
@@
where $H_{\theta}\mathcal{L}$ is the Hessian of $\mathcal{L}$ with respect to $\theta$. With the gradient computed so, we can run projected gradient ascent to approximate the attack.
Recall that in our threat model, we do not assume knowledge of the random seed of learning, even if we know the model architecture. To deal with this, we run the gradient-based algorithm from $K$ random neural network initializations. For each such initialization $k$, we obtain an approximately optimal attack $\delta^{k}$. We then use these in a kind of ensemble to compute the final poisoning attack as follows. First, compute $\delta=(\sum_{k}\delta^{k})/K$. Second, we choose the top $B$ $\delta_{i}$s in terms of the value of $|\delta_{i}|$ obtained in this fashion, and flip the associated labels $o_{i}$. As an ablation, we also consider an approach proposed by Koh and Liang (2017) in which they pre-train the neural network to initial gradient descent; in our experiments, however, we find that our approach tends to outperform the pre-trained variant. Algorithm 1 provides the details of the proposed gradient-based algorithm for the white-box setting.
@@
An important technical challenge with this gradient-based approach is that it becomes impractical in high dimensions because of the size of the Hessian combined with the inverse in the implicit gradient calculation. We consider three approaches for dealing with this. The first is similar to Koh and Liang (2017), and makes use of the conjugate gradient method. Specifically, instead of computing the inverse directly, we reframe implicit gradient computation as a solution to a linear system of equations $Ax=b$, where $A$ in our context corresponds to the Hessian of the loss. We use HVP (implicit Hessian-vector products) to approximate $Ax$ (Koh and Liang, 2017), and use Newton’s conjugate gradient to approximate the solution for $Ax=b$. However, we experimentally found that this approach performs quite poorly in our setting (either taking a very long time, or yielding poor efficacy). Our second approach is to first learn a model on clean data $D$, and then use a lower-dimensional embedding of the resulting reward model as the feature vector for outcomes in computing the attack (this is, of course, not accounting for the fact that the attack will also impact this embedding). Our third approach is to use PCA to significantly reduce input dimensionality, and apply the gradient-based approach with a proxy neural network architecture using the reduced feature vectors.
Our final consideration is the issue of black-box attacks, where the attacker does not know the model architecture. Note, however, that the use of dimensionality reduction such as PCA already entails learning a model that has a distinct architecture than the one being attacked (as the input dimension is different from original). More generally, our approach to such black-box attack problems follows the common pattern in prior literature where we attack a proxy model architecture, and subsequently evaluate transferability (Suciu et al., 2018).
@@
5.2. Rank-by-Distance Approaches

### v2

@@ 3 Threat Model
We consider the presence of a malicious actor (adversary) who can modify preference labels for pairs of outcomes in the dataset. Since these are inherently subject to human feedback, with no objective way to ground them (being expressions of human preferences), it is not possible in general to verify responses received against some ground truth. Consequently, even manual screening of such data cannot reliably identify poisoned instances, as these may simply be expression of unusual or unexpected preferences. Moreover, preferences of this kind are commonly obtained over the internet, the relative anonymity of which provides ample opportunity for malicious subversion of preference information. In particular, attackers can enter as annotators of the system, solicited, for example, using Amazon Mechanical Turk, Prolific, etc, or hired using other means (e.g., by using a third-party outsourcing service, such as Sama [64]), or can pay off a subset of such annotators.
3.1 Attacker Capabilities and Constraints
@@
We assume that the attacker can modify (flip) a subset of preference labels $o_{i}$ for corresponding pairs of input outcomes $(x_{i},y_{i})$, but not the feature vectors associated with these outcomes, as these are typically pre-specified as part of the preference elicitation process, and are consequently far more difficult to directly modify. We allow at most $B$ labels to be flipped. For example, if this is done through malware on a user’s device that submits fake pairwise comparison results on the user’s behalf, such attacks can affect only a relatively subset of users. Similarly, if malicious users interject themselves into the elicitation process (something that is difficult to prevent at scale, given the relative anonymity inherent in it), there is a limit to how many such malicious accounts can be created (or a limit on the cost an attacker can expend on using botnets for such a task). Finally, if attackers simply pay off regular hired annotators to submit malicious preferences, there is a limit on how much the adversary can spend.
@@
Let $\tilde{D}=\{(x_{i},y_{i},\tilde{o}_{i})\}$ be the dataset produced after the attacker manipulates (up to) $B$ labels in $D$, and let $\mathcal{L}(\tilde{D},\theta)$ be the loss function optimized by the learner over the poisoned dataset $\tilde{D}$. Let $\delta_{i}\in\{-1,0,+1\}$ denote the label perturbations by the attacker, so that $\tilde{o}_{i}=o_{i}+\delta$. The constraint is then $\sum_{i\in D}|\delta_{i}|\leq B$. We make explicit the dependence of poisoned data on $\delta$ using the notation $\tilde{D}(\delta)$.
3.2 Attacker Information
We consider two information settings vis-a-vis both the dataset and the model architecture: white-box and black-box. In the white-box model setting, the attacker is assumed to know the true model (e.g., neural network) architecture being trained. In the white-box data setting, we assume that the attacker knows the entire dataset $D$. In the black-box model setting, the attacker does not know the architecture. Notably, we do not assume knowledge of the random initialization in training. Finally, in the black-box data setting, we assume that the attacker knows, and can modify, only a subset of the full dataset $D$; in this setting, we systematically study the impact of partial knowledge of $D$ on the attack success rate.
3.3 Attacker Goals
@@ 3.3.1 Promotion Attacks
A key challenge in formalizing attacks that aim to promote a set $C^{T}$ of outcomes over others is that it is not evident how to identify these “others”. Typically, they concern unknown future comparisons, for example, in deciding whether a target product $x$ should be shown rather than some other product $x^{\prime}$ that was recently purchased by another user.
We approach this issue by viewing the attack goal from the perspective of learning theory [1]. Specifically, suppose that future products to which targets $x$ will be compared to will be generated according to some unknown distribution $\mathcal{P}$. We assume that the outcome pairs $(x,y)$ contained in the dataset $D$ are generated i.i.d. according to $\mathcal{P}$. This allows us to use alternatives in $D$ as a proxy in determining the objective to optimize in promotion attacks. The central theoretical question we ask in this section is whether the optimal solution to this proxy objective yields a provably approximately optimal promotion attack if comparisons are with respect to outcomes $x^{\prime}$ drawn randomly from $\mathcal{P}$, rather than from $D$. Next, we make this more precise.
We model the promotion attack as solving the following optimization problem:
@@
$\begin{split}&\max_{\delta}F(\delta)\equiv\sum_{c\in C^{T}}\Pr_{x\sim\mathcal{P}}\{R_{\theta}(c)\geq R_{\theta}(x)\}\\ &\mathrm{s.t.:}\quad\theta\in\arg\max_{\theta^{\prime}}\mathcal{L}(\tilde{D}(\delta),\theta^{\prime}).\end{split}$ (2)
@@
Notably, it is impossible to evaluate $\Pr_{x\sim\mathcal{P}}\{R_{\theta}(c)\geq R_{\theta}(x)\}$ since $\mathcal{P}$ is unknown. Moreover, even if we knew $\mathcal{P}$, computing this probability exactly can be computationally intractable. Practically, therefore, we would instead estimate it using a collection of sample outcomes $\{x_{i}\}_{i=1}^{N}$, which we assume are drawn from the unknown distribution $\mathcal{P}$ (a conventional assumption in learning theory, for example [1]). In practice, we would draw them, for example, from the dataset we aim to poison. This yields the following finite-sample approximation of Problem (2):
@@
$\begin{split}&\max_{\delta}\hat{F}(\delta)\equiv\sum_{c\in C^{T}}\sum_{i}\mathbf{1}_{R_{\theta}(c)\geq R_{\theta}(x_{i})}\\ &\mathrm{s.t.:}\quad\theta\in\arg\max_{\theta^{\prime}}\mathcal{L}(\tilde{D}(\delta),\theta^{\prime}),\end{split}$ (3)
where $\mathbf{1}_{R_{\theta}(c)\geq R_{\theta}(x_{i})}$ is 1 whenever the condition is true, and 0 otherwise. A natural question is whether this approximation yields a solution that approximates the true optimization problem that the attacker aims to solve—that is, Problem (2). We address this issue in Section 4.3 below.
3.3.2 Demotion Attacks
@@
Demotion attacks are a mirror image of promotion attacks: the attacker goal in this case is to demote a target set of candidates $C^{T}$ in terms of the comparisons with others induced by the learned reward model $R_{\theta}$. For example, a firm in a concentrated market would wish to prevent would want to demote the perceived preference of its competitor’s product, or a malicious actor would wish to prevent certain useful information from being shown by conversational AI (e.g., vaccine information in response to prompts about an infectious disease). We formalize it in precisely the same manner as we did for promotion attacks, with the goal now to minimize $F(\delta)$, rather maximize it. We can again use $\hat{F}(\delta)$ as the approximate objective.
3.3.3 Value Alignment and RLHF
An important, but far from sole, motivation for our consideration of attacks on reward model learning is RLHF. In this case, the learned—and, in our case, poisoned—reward models are then used downstream as rewards in a reinforcement learning loop (commonly, using PPO) to obtain policies that make decisions which are aligned with values represented by the reward model [62, 80, 78]. While not a primary part of our threat model, we evaluate the downstream effect of our attack on policies obtained using RLHF in Section 5 as well.
3.4 Stealth
@@
An important consideration in poisoning attacks is stealth, that is, ensuring that the fact of the attack is not immediately evident to observers. To this end, we use degradation in overall test accuracy as a measure of stealth: we view the attack as stealthy if test accuracy does not significantly degrade. We treat stealth as a post-hoc requirement on the attack, that is, we experimentally check that the attack does not have a significant accuracy degradation as a side-effect. An important factor in this context is that the semantics of accuracy in preference data are atypical, because preferences are inherently subjective, and people disagree. Consequently, accuracy variation can be easily attributable to a high disagreement rate, rather than a problem with the data itself, and accuracy degradation of over 10% need not reduce data credibility. However, once error is too high, stealth will indeed be undermined.
4 Attack Algorithms
There are three algorithmic challenges in developing algorithmic techniques for the promotion and demotion attacks. First, both Problem (2), and its finite-sample approximation (3) (as well as their demotion attack counterparts) entail solving a bi-level optimization problem. Second, our problem involves flipping a subset of preference comparison labels. Thus, in contrast to the majority of data poisoning approaches, including all gradient-based methods to date [5, 27, 31, 34, 111], which assume a continuous decision space (changing real-valued features), we face a combinatorial optimization problem over a discrete space of labels. Third, the objective function in the proxy formulation (3) is discontinuous and non-differentiable, precluding direct application of gradient-based methods.
We develop two classes of algorithmic approaches to address these challenges. First, we propose a novel projected gradient ascent algorithm for preference label poisoning attacks, and several variations that address its scalability challenges in high dimensions. The second class are greedy rank-by-distance approaches, in which the main variation is how we measure distance. We detail these next. Finally, we present a theoretical analysis that formally justifies our focus on the finite-sample objective (Problem (3)) as a proxy for the true adversarial goal (Problem (2)).
4.1 Gradient-Based Framework
@@
$U(\delta)=\sum_{c\in C^{T}}R_{\theta}(c)-\frac{|C^{T}|}{N}\sum_{i}R_{\theta}(x_{i}).$ (4)
Now, since the only dependence on the decision variables $\delta$ is via the parameters learned $\theta$, as long as $R_{\theta}$ are differentiable with respect to $\theta$, the objective will pose no issues. Next, we turn to the problem of discrete decision space $\delta$, where $\delta_{i}\in\{-1,0,+1\}$. Since the decision variables are discrete, we cannot directly use gradient-based methods for optimization. Our solution is a projected gradient descent (PGD) method in which we relax the variables $\delta_{i}$ to be in the interval $[-1,+1]$ (instead of being discrete), apply a sequence of gradient steps, and then project back to $\{-1,0,+1\}$. The projection step will choose the $B$ $\delta_{i}$s with the largest magnitude as the attack, setting the rest to zero, and then round the selected $\delta_{i}$s. Consequently, since the final step involves this projection, the approach is guaranteed to return a feasible $\delta$ vector.
@@
where we now make explicit the dependence of $\theta$ on $\delta$, and $\frac{d\theta(\delta)}{d\delta}$ is the matrix of derivatives $\frac{d\theta_{j}}{d\delta_{i}}$. While computing $\nabla_{\theta}R_{\theta}$ is straightforward, $\theta(\delta)$ is an implicit function of $\delta$ that arises due to the indirect dependence of learned parameters $\theta$ on $\delta$ as solutions to the optimization problem $\min_{\theta^{\prime}}\mathcal{L}(\tilde{D}(\delta),\theta^{\prime})$. Computing the implicit derivatives $\frac{d\theta(\delta)}{d\delta}$ is a common issue in gradient-based methods for poisoning attacks (see, e.g., [59]), and a common approach is to leverage the implicit function theorem [44, 59, 75]. Specifically, since MLE in our context is an unconstrained optimization problem, an optimal solution satisfies first-order conditions $\nabla_{\theta}\mathcal{L}(\tilde{D}(\delta),\theta)=0$. With mild assumptions detailed elsewhere [75], we compute the implicit derivative:
@@
where $H_{\theta}\mathcal{L}$ is the Hessian of $\mathcal{L}$ with respect to $\theta$. With the gradient computed so, we can run projected gradient ascent to approximate the attack. In our implementation, we use PyTorch autograd to compute the gradient and Hessian.
In our threat model we do not assume knowledge of the random seed used for training. To deal with this, we run the gradient-based algorithm from $K$ random neural network initializations. For each such initialization $k$, we obtain an approximately optimal attack $\delta^{k}$. We then use these in a kind of ensemble to compute the final poisoning attack as follows. First, compute $\delta=(\sum_{k}\delta^{k})/K$. Second, we choose the top $B$ $\delta_{i}$s in terms of the value of $|\delta_{i}|$ obtained in this fashion, and flip the associated labels $o_{i}$. As an ablation, we also consider an approach proposed by [44] in which they pre-train the neural network to initial gradient descent; in our experiments, however, we find that our approach tends to outperform the pre-trained variant. Algorithm 1 provides the details of the proposed gradient-based algorithm.
@@
An important technical challenge with this gradient-based approach is that it becomes impractical in high dimensions because of the size of the Hessian combined with the inverse in the implicit gradient calculation. We consider three approaches for dealing with this. The first is similar to [44], and makes use of the conjugate gradient method. Specifically, instead of computing the inverse directly, we reframe implicit gradient computation as a solution to a linear system of equations $Ax=b$, where $A$ in our context corresponds to the Hessian of the loss. We use HVP (implicit Hessian-vector products) to approximate $Ax$ [44], and use Newton’s conjugate gradient to approximate the solution for $Ax=b$. However, we experimentally found that this approach performs quite poorly in our setting (either taking a very long time, or yielding poor efficacy). Our second approach is to first learn a model on clean data $D$, and then use a lower-dimensional embedding of the resulting reward model as the feature vector for outcomes in computing the attack (this is, of course, not accounting for the fact that the attack will also impact this embedding). Our third approach is to use PCA to significantly reduce input dimensionality, and apply the gradient-based approach with a proxy neural network architecture using the reduced feature vectors.
Our final consideration is the issue of black-box attacks, where the attacker does not know the model architecture. Note, however, that the use of dimensionality reduction such as PCA already entails learning a model that has a distinct architecture than the one being attacked (as the input dimension is different from original). More generally, our approach to such black-box attack problems follows the common pattern in prior literature where we attack a proxy model architecture, and subsequently evaluate transferability [79].
@@
4.2 Rank-by-Distance Approaches